OpenStudy (anonymous):

If 6x2(6 x square)-5x-6=0,then what is the value of x? can any body help me?

OpenStudy (baldymcgee6):

\[6x^2 - 5x -6 = 0\] Like that?

OpenStudy (anonymous):

yes

OpenStudy (baldymcgee6):

Do you know how to factor something like this?

OpenStudy (anonymous):

no

OpenStudy (baldymcgee6):

Have you heard of the quadratic formula?

OpenStudy (anonymous):

yes ...but am looking for some easy approach to solve!!!

OpenStudy (baldymcgee6):

\[\frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a }\]

OpenStudy (baldymcgee6):

this is an easy method.. :)

OpenStudy (anonymous):

this question is from a GRE exam...my doubt how to do such a long equation within short time...do any body knows some easy approach?

OpenStudy (anonymous):

This can be factored.

OpenStudy (anonymous):

Here a = 6, b = -5, and c = -6. So you need to find all of the factors of (a * c) first. Can you do this?

OpenStudy (anonymous):

I am sorry i did't got your point